OpenStudy (kittiwitti1):

A 873 kg dragster, starting from rest, attains a speed of 26.3 m/s in 0.59 s. a) Find avg. acceleration: 45 m/s/s. FOUND b) Size of avg. force: 39000 N. FOUND c) Assume the driver has a mass of 68 kg. What horizontal force does the seat exert on the driver? 3100 N. FOUND d) Is the driver's mass in part c) an inertial or gravitational mass? --- A 65 kg swimmer jumps off a 10 m tower. a) find the swimmer's velocity when hitting the water. b) The swimmer comes to a stop 2 m below the surface. Find the net force exerted by the water.

OpenStudy (anonymous):

ok start solving.. what u waiting for? :P

OpenStudy (kittiwitti1):

Wait! editing

OpenStudy (anonymous):

Have you seen kinetic and potential energy?

OpenStudy (kittiwitti1):

@Lessis: No, I haven't ... we just started this topic ...

OpenStudy (kittiwitti1):

I need letter c) for the first one, and I have 3 more problems until the second one. Kinetic is energy in motion, right?

OpenStudy (anonymous):

\[E_{K} = \frac{ 1 }{ 2 }mv^{2} \]

OpenStudy (kittiwitti1):

... I don't recognize that formula...

OpenStudy (anonymous):

That's Kinetic Energy.

OpenStudy (kittiwitti1):

Well, yes I know that, but we haven't learned it. So we shouldn't be using it, right?

OpenStudy (anonymous):

You can use it, but I'm pretty sure it's not recommended. Do you know your vertical throw equations?

OpenStudy (kittiwitti1):

Eh. What...? o.o

OpenStudy (kittiwitti1):

All I know are these:\[F = ma\] \[\Sigma Fy = Fa - W = ma\]

OpenStudy (kittiwitti1):

And the two about acceleration and velocity that you told me.

OpenStudy (anonymous):

You can calculate the velocity of the swimmer using the equation I gave you for constant acceleration and initial velocity cero.

OpenStudy (kittiwitti1):

cero?

OpenStudy (anonymous):

After you calculate the velocity with which the swimmer touches the water, you can calculate his deceleration, considering that he stopped after two meters. (You need to calculate the acceleration required to get to a complete stop, from the velocity you calculated, to a velocity of 0, in just two meters) After you get that acceleration, you can calculate the force by multiplying by mass.

OpenStudy (kittiwitti1):

Is that Spanish for zero?

OpenStudy (kittiwitti1):

I haven't gotten to that question yet

OpenStudy (anonymous):

... Perdón. xD Fue una ligera falla de ortografía dado a que estoy acostumbrado (malacostumbrado?) a escribir cero. :P

OpenStudy (kittiwitti1):

& I think I'd need a diagram... no me gusta espanol = n=

OpenStudy (kittiwitti1):

wait I have another question! o.o (see pulse)

OpenStudy (kittiwitti1):

.... nvm i put it here After a day of testing race cars, you decide to take your own 1550 kg car onto the test track. While moving down the track at 10 m/s you suddenly accelerate to 30 m/s in 10 s. What is the avg. net force you have applied to the car during the 10s interval?

OpenStudy (kittiwitti1):

... lol "dun dun dunn" ?? good pic I get it

OpenStudy (anonymous):

\[Average Acceleration = \frac{ \Delta V }{ t }=\frac{ V_{2} - V_{1} }{ t } = \frac{ 30 \frac{ m }{ s } - 10\frac{m}{ s } }{ 10 s }\] \[Average Force = mass*averageacceleration\]

OpenStudy (kittiwitti1):

wut

OpenStudy (kittiwitti1):

\[a_{avg} = \frac{ \Delta v }{ t } = \frac{ v _{2} - v _{1}}{ t } = \frac{ 30 m/s ^{2} - 10 m/s ^{2}}{ 10s }\] \[F _{avg} = m \times a _{avg}\]

OpenStudy (kittiwitti1):

that's my idea of it

OpenStudy (anonymous):

Yes. But the velocities are given in meter per second, not meter per second squared.

OpenStudy (kittiwitti1):

-.- Grr. I forgot it was velocity & not acceleration.

OpenStudy (kittiwitti1):

Does that equal 20 m then?

OpenStudy (kittiwitti1):

2 m/s/s*

OpenStudy (anonymous):

The difference in speed is of 20 meters per second, and the acceleration is of 2 meters per second squared, yes.

OpenStudy (kittiwitti1):

#thumbsup

OpenStudy (kittiwitti1):

A race car has a mass of 710 kg. It starts from rest and travels 40 m in 3.0 s. The car is uniformly accelerated during the entire time. What net force is applied to it? Answer is 6300 N.

OpenStudy (kittiwitti1):

I get the velocity\[v \approx 13.3 m/s\]

OpenStudy (kittiwitti1):

This becomes \[a _{avg} = \frac{ v }{ t } = \frac{ 13.3 }{ 3.0 } \approx 4.44 m/s ^{2}\]Now I plug this in\[F = ma = 710 \times 4.44 \approx 3100 N\] I got the answer, I think.

OpenStudy (kittiwitti1):

Except a is not average... Did I do it right? @Lessis

OpenStudy (kittiwitti1):

Wait.... wtf NO I didn't get the answer = n =

OpenStudy (anonymous):

You need an equation I don't remember how to get. D: \[x =v_{0}t + \frac{ 1 }{ 2 }at^{2}\]

OpenStudy (kittiwitti1):

it's 4.43 not 4.44

OpenStudy (kittiwitti1):

... darn I hate those equations! Okay: \[x = v _{0}t + \frac{ 1 }{ 2 }at\]becomes\[40 = 0 + \frac{ 1 }{ 2 }a \left( 3.0 \right)\]which translates into\[80 = 3.0a\]and that makes\[a \approx 26.67\]which could be 27 or 26.7 depending on sig figs

OpenStudy (anonymous):

Time is squared.

OpenStudy (kittiwitti1):

m/s/s since I forgot. ... Crap. \[80 = 9.0a\]\[a \approx 8.89\] depending on sig figs, could be 8.9 or 9

OpenStudy (kittiwitti1):

\[8.9 m/s ^{2}\] it is. \[F = ma = 710 \times 8.9 \approx 6300 N\]

OpenStudy (anonymous):

Correct!

OpenStudy (kittiwitti1):

Thanks! :) A force of -9000 N is used to stop a 1500 kg car at 20 m/s. What braking distance is needed to bring the car to a halt?

OpenStudy (anonymous):

For the swimming problem, you need the same equation to calculate the force exerted by the water, by the way.

OpenStudy (kittiwitti1):

Ok, I gotcha. :p Do I still use F= ma?

OpenStudy (kittiwitti1):

For the last one I posted*

OpenStudy (kittiwitti1):

\[F =ma\]\[-9000 = 1500a\]\[a = -6\]

OpenStudy (anonymous):

Yes.

OpenStudy (kittiwitti1):

I'm guessing I find time and then use the equation to find \[\Delta x\]

OpenStudy (anonymous):

Use \[v = v_{0} + at\] To calculate the time needed for a complete stop. Then use \[x = v_{0}t + \frac{ 1 }{ 2 }at^{2} \] To calculate the distance.

OpenStudy (kittiwitti1):

Okay. \[v = v _{0} + \frac{ 1 }{ 2 }at ^{2}\]

OpenStudy (kittiwitti1):

That makes\[20 = 0 + \frac{ 1 }{ 2 }\left( -6 \right)t ^{2}\] am I right

OpenStudy (anonymous):

Wrong equation. You typed v instead of x. And V0 is your starting speed, which in this case is 20 m/s.

OpenStudy (kittiwitti1):

Wait.... I did the wrong equation completely #lolwhat

OpenStudy (anonymous):

\[v = v_{0} + at\] \[0 = 20 \frac{ m }{ s }+ at\]

OpenStudy (kittiwitti1):

\[20 = 0 + -6t\]

OpenStudy (kittiwitti1):

.-. oh right. \[0 = 20 m/s - 6t\]which is\[t = \frac{ -20 }{ -6 } \approx 3.3 s\]

OpenStudy (anonymous):

\[x = v_{0}t + \frac{ 1 }{ 2 }at^{2}\] x = (20*3.3) + 0.5*(-6)*(3.3)*(3.3)

OpenStudy (kittiwitti1):

\[x \approx 71 m\]wrong,it's 33 -.-

OpenStudy (kittiwitti1):

i messed up on the calculations

OpenStudy (kittiwitti1):

somewhere.... even though i use a calculator...

OpenStudy (kittiwitti1):

oh i left out the -6

OpenStudy (kittiwitti1):

ok time for swimmer problem

OpenStudy (kittiwitti1):

i got the pic

OpenStudy (kittiwitti1):

.-. i need time to solve///

OpenStudy (kittiwitti1):

I mean, I need t to solve

OpenStudy (anonymous):

\[x = v_{0}t + \frac{ 1 }{ 2 }at^{2}\]

OpenStudy (kittiwitti1):

well yes

OpenStudy (anonymous):

In this case, x is -10 meters. (You've fallen ten meters from the top of the pool) V0 is 0. (The speed at which you start, when you're at the top) a is g, which is equal to 9.8 m/s.

OpenStudy (anonymous):

a is also negative, considering how you're falling, by the way.

OpenStudy (kittiwitti1):

ok i got approx. 1.43

OpenStudy (anonymous):

That is your time, yes. Now, considering how the acceleration is constant use: \[a = \frac{ v-v_{0} }{ t } = \frac{ v }{ t }\] To calculate the speed.

OpenStudy (anonymous):

Remember that a is equal to g that is equal to -9.8 meters per second squared.

OpenStudy (kittiwitti1):

ok

OpenStudy (kittiwitti1):

9.8 = v/1.43?

OpenStudy (kittiwitti1):

I get 14.014

OpenStudy (anonymous):

Yes. Just remember that a was negative, so v must be negative. The negative velocity means that you're falling.

OpenStudy (kittiwitti1):

Vf=14 then?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now, you know that after touching the pool, he falls for two meters and then goes to a complete stop. So you have the next equation for the average acceleration in the water. \[a = \frac{ \Delta V }{ t }= \frac{ V_{2} - V_{1} }{ t } = \frac{ -V_{1} }{ t }\] Where v1 is the speed you just calculated, -14 m/s. So. \[a = -\frac{ -14 m/s }{ t } = \frac{ 14 m/s }{ t }\]

OpenStudy (anonymous):

You also know that: \[x = v_{0}t + \frac{ 1 }{ 2 }at^{2}\] Where V0 is your -14 m/s and x is -2 m.

OpenStudy (anonymous):

\[-2m = (-14m/s) t + \frac{ 1 }{ 2 }at^{2}\]

OpenStudy (anonymous):

So, now you have two equations and two variables (a and t). So you have to solve the system.

OpenStudy (kittiwitti1):

o.o I come back to see this..

OpenStudy (kittiwitti1):

wait

OpenStudy (kittiwitti1):

ok I got it, thanks @Lessis

OpenStudy (kittiwitti1):

#offline